OpenStudy (anonymous):

need to make sure my answer is right Mary has 27 pictures. She gives 1/3 of them to her sister barb and 2/3 to her couison Sara How many pictures does barb get 4 pictures. How mwny pictures does sara get? 5 how many pictures does mary keep herself 18 is that right

OpenStudy (dumbcow):

1/3 * 27 = 27/3 = 9 so she gives 9 to barb 2/3 *27 = 54/3 = 18 so she gives 18 to sara 18+9 =27 so there are none left over